OpenStudy (anonymous):

Help with the following: Find an equation of the line passing through each pair of point. Write equation in the form Ax=By=C

OpenStudy (bahrom7893):

what are the points?

OpenStudy (anonymous):

oh sorry, there (-1,3) and (-2,-5)

OpenStudy (bahrom7893):

And I assume u meant Ax + By = C

OpenStudy (anonymous):

yes, sorry

OpenStudy (bahrom7893):

so first we find the slope: M = (-5-3)/(-2+1) = -8/-1 = 8

OpenStudy (anonymous):

Wait can you still stay on mines lemme jus re do it and check to see if i get it rite..plz

OpenStudy (bahrom7893):

okay

OpenStudy (anonymous):

Where did you get -2+1, is'nt it suppose to be -2-1?

OpenStudy (bahrom7893):

-2 - (-1) = -2 + 1

OpenStudy (anonymous):

oh okay, wait..;plzz i really need help

OpenStudy (bahrom7893):

i will.. just yell at me in chat if im not here

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

okay i got the slope which i 8.. Then how to i put it in the ax+bx=c form

OpenStudy (bahrom7893):

okay now i would do this: y-y1 = M(x-x1) y-3 = 8(x+1)

OpenStudy (bahrom7893):

then simplify:

OpenStudy (anonymous):

SO its y-3=8x+8

OpenStudy (anonymous):

How do i simplify

OpenStudy (bahrom7893):

y -3 = 8x + 8 -8x + y = 8+3 -2x + y = 11

OpenStudy (bahrom7893):

So this is in form: Ax + By = C, where A = -2; B = 1 and C = 11

OpenStudy (anonymous):

How did you simplify??? And i keep getting -11=8x-y

OpenStudy (bahrom7893):

y -3 = 8x + 8 subtract 8x and add 3 from and to both sides: y - 8x - 3 + 3 = 8x - 8x + 3 + 8

OpenStudy (bahrom7893):

woops sorry that 2 was supposed to be an 8

OpenStudy (bahrom7893):

y - 8x = 11

OpenStudy (anonymous):

in the back of my book it gives me 8x-y=-11

OpenStudy (bahrom7893):

its the same thing, just multiply thru by negative 1... both of us are right

OpenStudy (anonymous):

so either way 8x-y=-11 and y-8x=11 are correct?